OpenStudy (anonymous):

If g(x) = 3(x − 2), find the value of x if g(x) = 6. 12 2.6 8 4

OpenStudy (4everaddicted2anime):

The answer is 4

OpenStudy (anonymous):

substitute 6 for x what do you get?

rvc (rvc):

instead of g(x) substitute 6

OpenStudy (anonymous):

If g(x) = x2 + 3, find g(4). 11 19 16 8

rvc (rvc):

6=3(x-2)

OpenStudy (anonymous):

expand the right side Then add both side by 6 Lastly divide both sides by 3

OpenStudy (4everaddicted2anime):

For the second question: Substitute 4 for any of the x values. (4)2+3 Multiply 4 and 2 8+3 Finally add to get 11

OpenStudy (anonymous):

For the function f(x) = x + 5, what is the ordered pair for the point on the graph when x = 4w? (4w, x + 5) (4w, 4w + 5) (x, x + 5) (x, 4w + 5)

OpenStudy (4everaddicted2anime):

The question gives you the x coordinate so now all you have to do is find the y coordinate. X=4w Substitute the x value into the equation to get the y coordinate. F(x)=4w+5 F(x) can be replaced by y. Y=4w+5 Final answer would be (4w,4w+5)

OpenStudy (anonymous):

Let f(p) be the average number of days a house stays on the market before being sold for price p in $1,000s. Which statement best describes the meaning of f(150)? This is the average number of days the house stayed on the market before being sold for $150,000. The house sold on the market for $150,000 and stayed on the market for an average of 150 days before being sold. The house sold for $150,000. The house stayed on the market for an average of 150 days before being sold.